(6.1) Find the values of a, b and c such the matrix below is skew symmetric.


0 0 d

0 2a − 3b + c 3a − 5b + 5c

2 0 5a − 8b + 6c


(6.2) Give an example of a skew symmetric matrix.

(6.3) Prove that A2 is symmetric whenever A is skewsymmetric.

(6.4) Determine an expression for det(A) in terms of det(AT) if A is a square skewsymmetric.

(6.5) Assume that A has an odd number of rows and also an odd number of columns. In this particular case, show that det(·) is an odd function.





1
Expert's answer
2021-05-17T13:41:02-0400

1) Given matrix is skew symmetric matrix-

    A'=-A


 "\begin{bmatrix}0&0&2\\0&2a-3b+c&0\\d&3a-5b+5c& 5a-8b+6c\end{bmatrix}=-\begin{bmatrix}0&0&d\\0&2a-3b+c&3a-5b+5c\\2&0&5a-8b+6c\end{bmatrix}"


On comparing we get-

"d=2, \\

2a-3b+c=-2a+3b-c"


"\Rightarrow 2a-3b+c=0
\\
\Rightarrow 3a-5b+5c=0
\\
\Rightarrow -5a+8b-6c=0"


On solving above equations we have-


a=0,b=0 and c=0


(2) Skew symmetric matrix is-


"A=\begin{bmatrix} 0&1&-2\\-1&0&3\\2&-3&0\end{bmatrix}"


(3) Let B=A^2


   if "A=A^T\Rightarrow A.A^T=A.A=A^2=B"

   "B^T=(A^TA)^T=(A^T)(A^T)^T=A^T.A=A.A=A^2=B"


  B is symmetric.

  Therefore "A^2" is symmetric.


(4)


"det(A^T)=|A^T|
\\[9pt]
 |-A|=(-1)^n|A|=-|A|
\\[9pt]
 det(A^T)=-det(A)
\\"


(5) Given that A matrix is a square matrix

 let n=3 be the order


"|-A|=(-1)^3|A|=-|A|


\\[9pt]
 |-A|=-|A|"


det() is an odd function